# Reinforcement Learning for Fine-tuning Text-to-speech Diffusion Models

### v1

@@ Abstract
Recent advancements in generative models have sparked significant interest within the machine learning community. Particularly, diffusion models have demonstrated remarkable capabilities in synthesizing images and speech. Studies such as those by Lee et al. [19], Black et al. [4], Wang et al. [37], and Fan et al. [8] illustrate that Reinforcement Learning with Human Feedback (RLHF) can enhance diffusion models for image synthesis. However, due to architectural differences between these models and those employed in speech synthesis, it remains uncertain whether RLHF could similarly benefit speech synthesis models. In this paper, we explore the practical application of RLHF to diffusion-based text-to-speech synthesis, leveraging the mean opinion score (MOS) as predicted by UTokyo-SaruLab MOS prediction system [29] as a proxy loss. We introduce diffusion model loss-guided RL policy optimization (DLPO) and compare it against other RLHF approaches, employing the NISQA speech quality and naturalness assessment model [21] and human preference experiments for further evaluation. Our results show that RLHF can enhance diffusion-based text-to-speech synthesis models, and, moreover, DLPO can better improve diffusion models in generating natural and high quality speech audios.
@@ 1 Introduction
Diffusion probabilistic models, initially introduced by Sohl-Dickstein et al. [33], have rapidly become the predominant method for generative modeling in continuous domains. Valued for their capability to model complex, high-dimensional distributions, these models are widely used in fields like image and video synthesis. Notable applications include Stable Diffusion [28] and DALL-E 3 [32], known for generating high-quality images from text descriptions. Despite these advancements, significant challenges persist where large-scale text-to-image models struggle to produce images that accurately correspond to text prompts. As highlighted by Lee et al. [19] and Fan et al. [8], current text-to-image models face challenges in composing multiple objects [9, 10, 26] and in generating objects with specific colors and quantities [13, 19]. Black et al. [4], Fan et al. [8], and Lee et al. [18] propose reinforcement learning (RL) approaches for fine-tuning diffusion models which can enhance their ability to align generated images with input texts.
@@
Figure 1: The steps of online RL for fine-tuning text-to-speech diffusion models: (1) Audios are sampled from the pretrained TTS model, (2) The reward model predicts a reward for the each sampled audio from step, (3) The text-to-image model is updated to maximize the reward.
In this study, we explore the application of online RL to fine-tune TTS diffusion models (Figure 1). Our evaluation focuses on the performance of TTS model fine-tuning using reward-weighted regression (RWR) [19], denoising diffusion policy optimization (DDPO) [4], diffusion policy optimization with KL regularization (DPOK)[8] and diffusion policy optimization with a KL-shaped reward [1], which we refer to as KLinR in this paper. We also introduce and assess diffusion loss-guided policy optimization (DLPO) where the reward is shaped by the diffusion model’s loss. In our experiments, we fine-tune WaveGrad 2, a non-autoregressive generative model for TTS synthesis [6], using reward predictions from the UTokyo-SaruLab mean opinion score (MOS) prediction system [29]. We find that RWR, DDPO, DPOK, and KLinR do not improve TTS models as they do for text-to-image models; because they cannot successfully control the magnitude of deviations from the original model. However, DLPO does, improving the sound quality and naturalness of the generated speech.
@@
    We are the first to apply reinforcement learning (RL) to fine-tune TTS diffusion models.
@@
    We introduce diffusion loss-guided policy optimization (DLPO). Unlike other RL methods, DLPO aligning with the training procedure of TTS diffusion models by incorporating the original diffusion model loss as a penalty in the reward function to effectively prevent model deviation and fine-tune TTS models.
@@
    We further investigate the impact of inference steps on TTS diffusion model fine-tuning by comparing a Monte Carlo approximation approach to fixed inference step sampling. We conduct a human experiment to evaluate the speech quality of DLPO-generated audio.
2 Related Work
@@
However, unlike text-to-image diffusion models, TTS diffusion models face the challenge of high temporal resolution. Audio input for TTS is a one-dimensional signal with a high sample rate; for example, one second of audio at 24,600 Hz consists of 24,600 samples. This high dimensionality requires managing thousands of data points, necessary to capture the nuances of human speech for natural-sounding audio. In contrast, a typical 256x256 image has 65,536 pixels, fewer than the samples in one second of high-fidelity audio. To handle and process the vast number of audio samples without compromising speed or accuracy, diffusion/denoising steps of TTS models are usually larger than that for text-to-image diffusion models, requiring memory optimization strategies during training. To address this, some TTS models only include part of the input sequence in training, which introduces a small mismatch between training and inference.
Reinforcement learning from human feedback.
Human assessments have been crucial in guiding model learning across various tasks, from behavior [18] and language modeling [3, 24] to text-to-image alignment [4, 8, 19]. This process typically involves collecting feedback at scale by having annotators compare outputs, and subsequently training a reward model to generate rewards aligned with human feedback. These reward models enhance the quality of text-to-image models through fine-tuning pre-trained generative models. Beyond these applications, human feedback has also optimized models in diverse settings including robotic control [7], game-playing [16], and more.
@@
Fan et al. [8] suggest this issue likely arises from fine-tuning on a fixed dataset generated by a pre-trained model. Black et al. [4] argue that the reward-weighted regression lacks theoretical grounding and only roughly approximates optimizing denoising diffusion with RL and they propose denoising diffusion policy optimization (DDPO), a policy gradient algorithm that outperforms reward-weighted regression methods. DDPO improves text-to-image diffusion models by targeting objectives like image compressibility and aesthetic appeal, and it enhances prompt-image alignment using feedback from a vision-language model, eliminating the need for additional data or human annotations. Fan et al. [8] also show that RL fine-tuning can surpass reward-weighted regression in optimizing rewards for text-to-image diffusion models. Additionally, they demonstrate that using KL regularization in RL methods helps address issues like image quality deterioration in fine-tuned models.
For fine-tuning TTS diffusion model, we develop an diffusion policy optimization shaped by diffusion model objective. This method outperforms RWR, DDPO, and DPOK and KLinR. We find that the latter methods, although reported to enhance text-to-image diffusion models, do not yield improvements in TTS model quality.
@@ 3.1 Text-to-Speech Diffusion Model
Figure 2: WaveGrad2 follows a Markov process where forward diffusion $q(x_{t+1}|x_{t})$ iteratively adds Gaussian noise to the signal starting from the waveform $y_{0}$. $q(x_{t+1}|x_{0})$ is the noise distribution used for training. The inference denoising process progressively removes noise, starting from Gaussian noise $x_{T}$. This figure is adapted from Chen et al. [5], Ho et al. [11].
Diffusion models gradually transform random noise into structured outputs, like speech, through a series of learned transformations. This involves two phases: the forward phase, where noise is incrementally added until only noise remains, and the reverse phase, where a model learns to reverse this noise addition to generate the desired output from pure noise.In this study, we use WaveGrad2 [6] as the pretrained TTS diffusion model for its ability to balance inference speed and sample quality by adjusting refinement steps and produce high-fidelity audio. Due to the lack of official code, we use an unofficial version from MINDs Lab (https://github.com/maum-ai/wavegrad2). WaveGrad2 is a non-autoregressive TTS model adapting the diffusion denoising probabilistic model (DDPM) from Ho et al. [11]. As shown in Figure 2, WaveGrad2 models the conditional distribution $p_{\theta}(x_{0}|c)$ where $x_{0}$ represents the waveform and $c$ the associated context. The distribution follows the reverse of a Markovian forward process $q(x_{t}|x_{t-1})$, which iteratively introduces noise to the data. Reversing the forward process can be accomplished by training a neural network $\mu_{\theta}(x_{t},c,t)$ with the following objective:
@@
where $\tilde{\epsilon}(x_{t},t)$ is the ground truth added noise for step $x_{t}$ and $\epsilon_{\theta}(x_{t},c,t)$ is the predicted noise for step $x_{t}$. As discussed in section 2, given the high waveform resolution in WaveGrad2 (24,000 samples per second), computing the loss across all waveform samples in an utterance is impractical due to significant computational and memory demands. Instead, after learning the representations on the whole input sequence, a small segment is sampled to synthesize the waveform. Random segments are sampled individually in each minibatch and the corresponding waveform segment is extracted based on the upsampling rate (300 in WaveGrad2 setup). WaveGrad2 employs a strategy to randomly select a specific time step $t$ and its associated noise, then calculates the loss by predicting the noise at this step. During inference, the entire input sequence is utilized, which introduces a small mismatch between training and inference. Additionally, to generate and process the vast number of audio sampled at audio at 24 kHz without compromising speed or accuracy, the number of diffusion/denoising steps is set to 1000 in WaveGrad2. This contrasts with text-to-image diffusion models from previous studies, which utilize the entire input sequence during both training and inference phases and involve 50 diffusion/denoising steps throughout their inference processes.
@@
Additionally, Chen et al. [5, 6] found that replacing the original $L^{2}$ distance metric with $L^{1}$ improves training stability. However, the unofficial WaveGrad2 implementation by MINDs Lab uses the original $L^{2}$ metric.
@@ 3.2 Reward Model
We use the UTokyo-SaruLab mean opinion score (UTMOS) prediction system [29] to predict the speech quality of generated audios from Wavegrad2. Mean opinion score (MOS) is a subjective scoring system that allows human evaluators to rate the perceived quality of synthesized speech on a scale from 1 to 5, which is one of the most commonly employed evaluation methods for TTS system [35]. UTMOS is trained to capture nuanced audio features that are indicative of human judgments of speech quality, providing an accurate MOS prediction without the need for extensive labeled data. Thus, we use UTMOS as the reward model for fine-tuning Wavegrad2.
@@ 4.3 Denoising Diffusion Policy Optimization
RWR relies on an approximate log-likelihood by disregarding the sequential aspect of the denoising process and only utilizing the final samples $x_{0}$. Black et al. [4] propose the denoising diffusion policy optimization to directly optimize $\mathcal{J}_{DDRL}$ using the score function policy gradient estimator, also known as REINFORCE [38, 22]. DDPO alternately collects denoising trajectories ${x_{T},x_{T-1},...,x_{0}}$ via sampling and updates parameters via gradient descent:
@@
In the training process of WaveGrad2, as described in subsection 3.1, it is computationally and memory-intensive to compute the loss for every waveform sample in a speech segment. Therefore, WaveGrad2 employs a strategy where it randomly selects a specific time step $t$ along with its associated noise. The loss is then calculated by predicting the noise at this selected time step. We can adapt DDPO to the WaveGrad2 setting as follows by leveraging a Monte Carlo approximation of the $\mathcal{J}_{\mathrm{DDRL}}$ objective:
$\nabla_{\theta}\mathcal{J}_{\mathrm{DDRL}}(\theta)=\mathbb{E}_{c\sim p(c)}\mathbb{E}_{p_{\theta}(x_{0:T}|c)}\mathbb{E}_{t\sim\mathcal{U}\{1,T\}}\left[\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)r(x_{0},c)\right]$ (8)
@@ 4.4 Diffusion Policy Optimization with a KL-shaped Reward
Fan et al. [8] demonstrate that adding KL between the fine-tuned and pre-trained models for the final image as a regularizer $KL(p_{\theta}(x_{0}|z)\|p_{\mathrm{pre}}(x_{0}|z)$ to the objective function helps to mitigate overfitting of the diffusion models to the reward and prevents excessively diminishing the "skill" of the original diffusion model. As discussed in subsection 3.1, $p_{\theta}(x_{0}|z)$ is calculated as a variational bound, Fan et al. [8] propose to add an upper-bound of this KL-term to the objective function $\mathcal{J}_{DDRL}(\theta)$, which they call DPOK. They find that DPOK outperforms reward weighted regression in fine-tuning text-to-image diffusion models:
@@
As DDPO, DPOK cannot be applied directly to WaveGrad2 for memory reasons, but we can adapt it using a Monte Carlo approximation:
$\mathbb{E}_{c\sim p(c)}\mathbb{E}_{p_{\theta}(x_{0:T}|c)}\mathbb{E}_{t\sim\mathcal{U}\{1,T\}}\left[-\alpha r(x_{0},c)\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)+\beta\nabla_{\theta}\log\frac{p_{\theta}(x_{t-1}|x_{t},c)}{p_{\mathrm{pre}}(x_{t-1}|x_{t},c)}\right]$ (10)
@@
where $\beta$ is the KL weight. We can adapt it to WaveGrad2 similarly:
$\nabla_{\theta}\mathcal{J}_{DDRL}(\theta)=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}c\sim p(c)\\ p_{\theta}(x_{0:T}|c)\\ t\sim\mathcal{U}\{1,T\}\end{subarray}}\Bigg{[}-\Bigg{(}\alpha r(x_{0},c)-\\ \beta\nabla_{\theta}\log\frac{p_{\theta}(x_{t-1}|x_{t},c)}{p_{\mathrm{pre}}(x_{t-1}|x_{t},c)}\Bigg{)}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)\Bigg{]}$
We evaluate both approaches with KL for fine-tuning WaveGrad2; for KLinR approach, we follow Fan et al. [8] and calculate the KL upper-bound. The algorithms are shown in Appendix A and Appendix B.
@@ 4.5 Diffusion Model Loss-Guided Policy Optimization
$\mathbb{E}_{c\sim p(c)}\mathbb{E}_{t\sim\mathcal{U}\{1,T\}}\mathbb{E}_{p_{\theta}(x_{0:T}|c)}\left[-\alpha r(x_{0},c)-\beta\|\tilde{\epsilon}(x_{t},t)-\epsilon_{\theta}(x_{t},c,t)\|_{2}\right]$ (12)
@@
$\mathbb{E}_{c\sim p(c)}\mathbb{E}_{t\sim\mathcal{U}\{0,T\}}\mathbb{E}_{p_{\theta}(x_{1:T}|c)}\left[-\left(\alpha r(x_{0},c)-\beta\nabla_{\theta}\|\tilde{\epsilon}(x_{t},t)-\epsilon_{\theta}(x_{t},c,t)\|_{2}\right)\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)\right]$ (13)
where we follow Ahmadian et al. [1] and add the diffusion model objective to the reward function as a penalty. The pseudo-code of our algorithm, which we refer to as DLPO, is summarized in Algorithm Algorithm 1. This algorithm aligning with the training procedure of TTS diffusion models by incorporating the original diffusion model objective $\beta\|\tilde{\epsilon}(x_{t},t)-\epsilon_{\theta}(x_{t},c,t)\|_{2}$ as a penalty in the reward function effectively prevent model deviation.

### v2

# DLPO: Diffusion Model Loss-Guided Reinforcement Learning for Fine-Tuning Text-to-Speech Diffusion Models

@@ Abstract
Recent advancements in generative models have sparked significant interest within the machine learning community. Particularly, diffusion models have demonstrated remarkable capabilities in synthesizing images and speech. Studies such as those by Lee et al. [20], Black et al. [3], Wang et al. [47], and Fan et al. [7] illustrate that Reinforcement Learning with Human Feedback (RLHF) can enhance diffusion models for image synthesis. However, due to architectural differences between these models and those employed in speech synthesis, it remains uncertain whether RLHF could similarly benefit speech synthesis models. In this paper, we explore the practical application of RLHF to diffusion-based text-to-speech with long-chain diffusion directly on waveform data, leveraging the mean opinion score (MOS) as predicted by UTokyo-SaruLab MOS prediction system [36] as a proxy loss. We introduce diffusion model loss-guided RL policy optimization (DLPO) and compare it against other RLHF approaches, employing the NISQA speech quality and naturalness assessment model [24] and human preference experiments for further evaluation. Our results show that RLHF can enhance diffusion-based text-to-speech synthesis models, and, moreover, DLPO can better improve diffusion models in generating natural and high-quality speech audio.
@@ 1 Introduction
Diffusion probabilistic models, initially introduced by Sohl-Dickstein et al. [41], have rapidly become the predominant method for generative modeling in continuous domains. Valued for their capability to model complex, high-dimensional distributions, these models are widely used in fields like image and video synthesis. Notable applications include Stable Diffusion [35] and DALL-E 3 [40], known for generating high-quality images from text descriptions. Despite these advancements, significant challenges persist where large-scale text-to-image models struggle to produce images that accurately correspond to text prompts. As highlighted by Lee et al. [20] and Fan et al. [7], current text-to-image models face challenges in composing multiple objects [8, 9, 30] and in generating objects with specific colors and quantities [13, 20]. Black et al. [3], Fan et al. [7], Lee et al. [19] propose reinforcement learning (RL) approaches for fine-tuning diffusion models which can enhance their ability to align generated images with input texts.
@@
In this study, we explore the application of online RL to fine-tune TTS diffusion models (Appendix A), especially whether RL techniques can similarly improve the naturalness and sound quality of diffusion TTS models that are directly trained on waveforms and assess the impact of RL on waveform generation in diffusion models. While basic policy gradient methods simply optimize the reward function, these can severely disrupt the initial model due to over-optimization. Therefore, commonly used RL methods control deviation from the original policy, often using estimates of the KL divergence between the learned policy and the original. In particular, we compare the performance of TTS model fine-tuning using reward-weighted regression (RWR) [20], denoising diffusion policy optimization (DDPO) [3], diffusion policy optimization with KL regularization (DPOK) [7] and diffusion policy optimization with a KL-shaped reward [1], which we refer to as KLinR in this paper. Both DPOK and KLinR use a diffusion gradient regularized KL. We also introduce and assess diffusion loss-guided policy optimization (DLPO) where the reward is shaped by the diffusion model’s loss. In our experiments, we fine-tune WaveGrad 2, a non-autoregressive generative model for TTS synthesis [5], using reward predictions from the UTokyo-SaruLab mean opinion score (MOS) prediction system [36]. We find that RWR and DDPO do not improve TTS models as they do for text-to-image models, because they cannot successfully control the magnitude of deviations from the original model. However, DPOK, KLinR, and DLPO do, improving the sound quality and naturalness of the generated speech, and DLPO outperforms DPOK and KLinR. Demo audios are presented in this website https://demopagea.github.io/DLPO_demo/
@@
    We are the first to apply reinforcement learning (RL) to improve the speech quality of TTS diffusion models.
@@
    We introduce diffusion loss-guided policy optimization (DLPO). Unlike other RL methods, DLPO aligns with the training procedure of TTS diffusion models by incorporating the original diffusion model loss as a penalty in the reward function to effectively prevent model deviation and fine-tune TTS models.
@@
    We further investigate the impact of diffusion gradients on fine-tuning TTS diffusion models and the effect of different denoising steps. We conduct a human experiment to evaluate the speech quality of DLPO-generated audio.
2 Related Works
@@
However, unlike text-to-image diffusion models, TTS diffusion models face the challenge of high temporal resolution. Audio input for TTS is a one-dimensional signal with a high sample rate; for example, one second of audio at 24,600 Hz consists of 24,600 samples. This high dimensionality requires managing thousands of data points, necessary to capture the nuances of human speech for natural-sounding audio. In contrast, a typical 256x256 image has 65,536 pixels, fewer than the samples in one second of high-fidelity audio. Handling and processing the large volume of audio samples while maintaining both speed and high speech quality typically necessitates diffusion and denoising steps in TTS models that are larger than those required for text-to-image diffusion models. This demands effective memory optimization strategies during training. Most prior works in this area try to address the problem of processing large volume of audio samples by preprocessing the high-dimensional waveform into Mel-spectrograms—two-dimensional images of time and frequency which simplify the problem to image synthesis, such as Grad-TTS [31, 34, 21, 6, 16, 44, 10]. Other models like NaturalSpeech 2 [39], SimpleTTS [23], DiTTo-TTS [18] and VALL-E [45] employ neural audio codecs to convert high-dimensional raw waveforms into quantized latent vectors. Then a vocoder is introduced to predict the audio from these intermediate features. Appendix G shows the state of the art for recent TTS models. WaveGrad2 performs comparably to GradTTS, which uses a vocoder, and outperforms E3 TTS and SimpleTTS. Other recent models are trained on massive proprietary datasets, which makes it difficult to compare their performance with earlier work or to use them as a starting point for basic research. While vocoder models are inherently more efficient and easier to tune than models based on the raw waveform due to using lower-dimensional inputs, Gao (2023) argues that they can be less domain-general due to the possibility of cascading errors from the preprocessor. Rather than trying to reduce TTS to vision by changing the inputs, we therefore select WaveGrad2 as our baseline model and try to modify RL techniques to optimize it.
@@
Fan et al. [7] suggest this issue likely arises from fine-tuning on a fixed dataset generated by a pre-trained model. Black et al. [3] argue that the reward-weighted regression lacks theoretical grounding and only roughly approximates optimizing denoising diffusion with RL and they propose denoising diffusion policy optimization (DDPO), a policy gradient algorithm that outperforms reward-weighted regression methods. DDPO improves text-to-image diffusion models by targeting objectives like image compressibility and aesthetic appeal, and it enhances prompt-image alignment using feedback from a vision-language model, eliminating the need for additional data or human annotations. Fan et al. [7] also show that RL fine-tuning can surpass reward-weighted regression in optimizing rewards for text-to-image diffusion models. Additionally, they demonstrate that using a diffusion gradient regularized KL in RL methods helps address issues like image quality deterioration in fine-tuned models.
For fine-tuning TTS diffusion model, [26] explores various RL techniques to improve emotional expression in Grad-TTS [31], including Reward Weighted Regression (RWR) and Proximal Policy Optimization (PPO). They demonstrate that Grad-TTS can convey emotions more effectively by utilizing feedback from an emotion predictor model. However, the speech quality and naturalness of their demo audio remain insufficient. We propose a diffusion policy optimization method guided by the diffusion model objective, which surpasses RWR, DDPO, DPOK, and KLinR. While RWR and DDPO have been shown to enhance text-to-image diffusion models, we find they do not improve speech quality in TTS models. Both DPOK and KLinR utilize diffusion gradient regularized KL, which has shown some improvements in speech quality, but our approach DLPO, which directly use diffusion loss as reward, achieves the best results.
@@ 3.1 Text-to-Speech Diffusion Model
In this study, we use the PyTorch implementation of WaveGrad2 provided by MINDs Lab (https://github.com/maum-ai/wavegrad2) as the pretrained TTS diffusion model. WaveGrad2 is a non-autoregressive TTS model adapting the diffusion denoising probabilistic model (DDPM) from Ho et al. [11]. WaveGrad2 models the conditional distribution $p_{\theta}(y_{0}|x)$ where $y_{0}$ represents the waveform and $x$ the associated context. The distribution follows the reverse of a Markovian forward process $q(y_{t}|y_{t-1})$, which iteratively introduces noise to the data.
Reversing the forward process can be accomplished by training a neural network $\mu_{\theta}(x_{t},c,t)$ with the following objective:
@@
where $\tilde{\epsilon}(x_{t},t)$ is the ground truth added noise for step $x_{t}$ and $\epsilon_{\theta}(x_{t},c,t)$ is the predicted noise for step $x_{t}$. More details are provided in Appendix B.
@@ 3.2 Reward Model
We use the UTokyo-SaruLab mean opinion score (UTMOS) prediction system [36] to predict the speech quality of generated audios from Wavegrad2. Mean opinion score (MOS) is a subjective scoring system that allows human evaluators to rate the perceived quality of synthesized speech on a scale from 1 to 5, which is one of the most commonly employed evaluation methods for TTS system [43]. UTMOS is trained to capture nuanced audio features that are indicative of human judgments of speech quality, providing an accurate MOS prediction without the need for extensive labeled data. Thus, we use UTMOS as the reward model for fine-tuning Wavegrad2. UTMOS is trained on datasets from the VoiceMOS Challenge 2022 [14], including 14 hours audio from male and female speakers with MOS ratings.
@@ 4.3 Denoising Diffusion Policy Optimization
RWR relies on an approximate log-likelihood by disregarding the sequential aspect of the denoising process and only using the final samples $x_{0}$. Black et al. [3] propose the denoising diffusion policy optimization to directly optimize $\mathcal{J}_{DDRL}$ using the score function policy gradient estimator, also known as REINFORCE [48, 25]. DDPO alternately collects denoising trajectories ${x_{T},x_{T-1},...,x_{0}}$ via sampling and updates parameters via gradient descent:
@@ 4.4 Diffusion Policy Optimization with a KL-shaped Reward
Fan et al. [7] demonstrate that adding KL between the fine-tuned and pre-trained models for the final image as a regularizer $KL(p_{\theta}(x_{0}|z)\|p_{\mathrm{pre}}(x_{0}|z)$ to the objective function helps to mitigate overfitting of the diffusion models to the reward and prevents excessively diminishing the "skill" of the original diffusion model. As discussed in subsection 3.1, $p_{\theta}(x_{0}|z)$ is calculated as a variational bound, Fan et al. [7] propose to add an upper-bound of this KL-term to the objective function $\mathcal{J}_{DDRL}(\theta)$, which they call DPOK. In their model implementation, this KL-term is computed as $\left[\|\epsilon_{\theta}(x_{t},c,t)-\epsilon_{pre}(x_{t},c,t)\|_{2}\right]$ following diffuson model objective. They find that DPOK outperforms reward weighted regression in fine-tuning text-to-image diffusion models:
@@
where $\beta$ is the KL weight. We evaluate both approaches with KL for fine-tuning WaveGrad2; for KLinR approach, we follow Fan et al. [7] and calculate the KL upper-bound. The algorithms are shown in Appendix C and Appendix D.
@@ 4.5 Diffusion Model Loss-Guided Policy Optimization
$\mathbb{E}_{c\sim p(c)}\mathbb{E}_{p_{\theta}(x_{0:T}|c)}\left[-\alpha r(x_{0},c)-\beta\|\tilde{\epsilon}(x_{t},t)-\epsilon_{\theta}(x_{t},c,t)\|_{2}\right]$ (10)
@@
$\mathbb{E}_{c\sim p(c)}\mathbb{E}_{p_{\theta}(x_{1:T}|c)}\left[-\left(\alpha r(x_{0},c)-\beta\nabla_{\theta}\|\tilde{\epsilon}(x_{t},t)-\epsilon_{\theta}(x_{t},c,t)\|_{2}\right)\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)\right]$ (11)
where we follow Ahmadian et al. [1] and add the diffusion model objective to the reward function as a penalty. The pseudocode of our algorithm, which we refer to as DLPO, is summarized in Algorithm 1. This algorithm aligning with the training procedure of TTS diffusion models by incorporating the original diffusion model objective $\beta\|\tilde{\epsilon}(x_{t},t)-\epsilon_{\theta}(x_{t},c,t)\|_{2}$ as a penalty in the reward function effectively prevents model deviation. More details and an overview figure are provided in Appendix A.